# Longteng Guo1,3, Jie Jiang1,3, Jing Liu1,3 1Zidongtaichu Foundation Model Research Center, CASIA 2State Key Laboratory of Multimodal Artificial Intelligence Systems, CASIA

### v1

@@ Abstract
Electroencephalogram (EEG) signals are pivotal in providing insights into spontaneous brain activity, highlighting their significant importance in neuroscience research. However, the exploration of versatile EEG models is constrained by diverse data formats, outdated pre-training paradigms, and limited transfer learning methods, only leading to specialist models on single dataset. In this paper, we introduce EEGPT, the first generalist EEG foundation model designed to address these challenges. First, we propose an electrode-wise modeling strategy that treats each electrode as a fundamental unit, enabling the integration of diverse EEG datasets collected from up to 138 electrodes, amassing 37.5M pre-training samples. Second, we develop the first autoregressive EEG pre-trained model, moving away from traditional masked autoencoder approaches to a next signal prediction task that better captures the sequential and temporal dependencies of EEG data. We also explore scaling laws with model up to 1.1B parameters — the largest in EEG research to date. Third, we introduce a multi-task transfer learning paradigm using a learnable electrode graph network that is shared across tasks, which for the first time confirms multi-task compatibility and synergy. As the first generalist EEG foundation model, EEGPT shows broad compatibility with various signal acquisition devices, subjects, and tasks. It supports up to 138 electrodes and any combination thereof as input. Furthermore, we simultaneously evaluate it on 5 distinct downstream tasks across 12 benchmarks. EEGPT consistently outperforms existing specialist models across all downstream tasks, with its effectiveness further validated through extensive ablation studies. This work sets a new direction for generalist EEG modeling, offering improved scalability, transferability, and adaptability for a wide range of EEG applications. Both the training code and model checkpoints will be publicly available.
Figure 1: EEGPT, as a generalist model, significantly outperforms dataset-specific specialist models across 12 benchmarks spanning 5 tasks. It strongly demonstrate the versatility and transferability.
1 Introduction
@@
Electroencephalogram (EEG), which captures spontaneous brain activity via electrograms (Biasiucci et al., 2019), could be conceptualized as the language of the brain. Through the analyses of EEG, valuable insights are derived for various applications, including but not limited to emotion recognition (Zhang et al., 2024), motor imagery classification (An et al., 2023), mental workload detection (Wang et al., 2024b) and sleep stage classification (Liang et al., 2023). This breadth of applications underscores the versatility and utility of EEG in neuroscientific research.
Research on EEG downstream tasks has been thriving, yet most studies share a notable characteristic: specialization. For instance, at the data level, a variety of proprietary data formats (Jia et al., 2020; Bashivan et al., 2015) and handcrafted feature extraction (Duan et al., 2013; Yan et al., 2023) techniques have been introduced to enhance the discriminability of domain-specific data. At the model level, various modules and structures are designed and trained for a specific task (Shao et al., 2023), dataset (Wang et al., 2024a), or even individual subjects (Gao et al., 2024). However, a generalist EEG foundation model is highly anticipated, as it offers broader applicability across various EEG tasks. Moreover, this model improves transfer learning by allowing knowledge from one task to enhance performance on others. Its design also demonstrates greater robustness to data and task variations, leading to better generalization in unseen scenarios.
Although extensive research in the fields of computer vision (CV) (Radford et al., 2021; Dosovitskiy, 2020; Bai et al., 2024) and natural language processing (NLP) (Radford et al., 2019; Achiam et al., 2023; Brown, 2020) has identified three key components for constructing generalist models—data, self-supervised pre-training, and transfer learning paradigms—EEG introduces its own unique and daunting challenges in each of these domains:
@@
Data Format. EEG data exhibit significant heterogeneity (Wang et al., 2024a; Saeed et al., 2021), characterized by a variety of systems (e.g., the 10-20 system) and equipment (e.g., Neuroscan) used in data collection. Furthermore, different datasets may employ a diverse number and combination of electrodes based on practical considerations. The inconsistency in data formats across different sources prevents their combined use in the same model for training, making it challenging to develop a generalist for EEG. Therefore, an efficient and scalable strategy for unifying these diverse EEG data format is extremely demanding.
Self-supervised Pre-training. Current studies (Yang et al., 2024b; Jiang et al., 2024; Yi et al., 2024) have uniformly employed techniques that mask parts of EEG signals and utilize a bidirectional attention mechanism (Vaswani, 2017) to reconstruct the masked data (i.e., mask autoencoder, MAE). However, they have inevitable limitations in capturing the sequential and temporal dependencies inherent in time-based data such as language and EEG. Given the gradual obsolescence of MAE architectures in NLP (Zhao et al., 2023; Minaee et al., 2024), it is essential for the EEG field, which shares similar temporal dynamics with language, to reconsider its current pre-training paradigms.
Transfer Learning. Current pre-trained EEG models are generally fine-tuned for specific datasets, resulting in specialists in narrow domains. However, in CV and NLP, many pre-trained models (Touvron et al., 2023; Yang et al., 2024a; Wang et al., 2023) have achieved remarkable generalizability through more adaptable and efficient knowledge transfer learning methods. These models support multiple tasks and promote beneficial synergies. Compared to models specialized in single tasks, they exhibit enhanced and broader capabilities, facilitating a more thorough utilization of pre-trained knowledge. However, advanced transfer learning method remains underexplored in the EEG field.
@@
In this paper, we propose EEGPT, a generalist EEG foundation model offering extensive versatility. Specifically, it seamlessly adapts and encodes signals collected by nearly all popular EEG acquisition devices. It accommodates signals from up to 138 electrodes, supporting various configurations and combinations. Moreover, EEGPT is capable of simultaneously processing and analyzing data from nearly all prevalent downstream tasks within a single model, and it is highly scalable to new tasks. The training recipe for EEGPT significantly diverges from previous paradigms, with its novelty encapsulated in five distinct "firsts":
@@
Figure 2: The versatility of EEGPT is reflected in the broad compatibility with subjects, signal acquisition devices, and tasks. EEG signals from various subject, using various device, and performing various task can be characterized and analyzed effectively within one model, and it exhibits remarkable scalability.
1) For data format, we propose the first electrode-wise modeling strategy. It deconstructs the signals electrode by electrode. Each electrode serves as a fundamental unit for subsequent training. Although the sets of electrodes differ across various datasets, this strategy consistently translates into an electrode-conditioned temporal modeling task. Leveraging this compatibility and unification, we extensively collect a total of 37.5M pre-training samples.
2) For self-supervised pre-training, we propose the first autoregressive EEG pre-trained model, seamlessly accommodating the sequential and temporal dependencies inherent in EEG data. Compared to MAE, the pre-trained model engages in a more intuitive yet challenging task of "next signal prediction". Based on 37.5M training samples, we split approximately 1B tokens to conduct pre-training across four scales (i.e., Base, Huge, Large, and Giant). To the best of our knowledge, It is the first exploration and validation of the scaling laws for autoregressive architectures in the EEG domain. Besides, EEGPT-Giant has achieved about 1.1B parameters, marking it as the first model in the EEG field to exceed the billion-parameter threshold.
@@
3) For transfer learning, we propose a learnable graph network, with electrodes as nodes, is concurrently shared across multiple tasks. Task-specific node activation patterns are adaptively determined by corresponding input data format. Leveraging the robust temporal representations learned from electrode-conditioned pre-training, the electrode graph serves as a spatial supplement by integrating information from multiple electrodes. The whole framework is designed with a progressive spatiotemporal decoupling. We collect data from 12 benchmarks for multi-task learning instead of traditional single-task fine-tuning. Interestingly, the tasks demonstrate a notable mutual enhancement. It establishes EEGPT as the first generalist EEG model for multi-task compatibility and synergism.
Based on these designs, our contributions are summarized as follows:
  • Electrode-wise Modeling Strategy. We introduce a novel electrode-wise strategy for EEG data integration, treating each electrode as a fundamental unit across various datasets. This approach enables uniform handling and scalability in data processing. Benefiting from this method, EEGPT can support up to 138 electrodes and their arbitrary combinations, offering flexibility and applicability far beyond existing models.
  • Autoregressive EEG Pre-trained Model: We introduce the first autoregressive pre-trained EEG model. Compared to traditional MAE techniques, it more naturally and efficiently captures the sequential and temporal dynamics inherent in EEG data. The scaling laws for data and model size in the autoregressive framework have been effectively validated.
@@
  • Multi-task Transfer Learning Paradigm: Building upon a learnable task-shared graph network, EEGPT is the first generalist model to exhibit confirmed multi-task compatibility and synergism. Significant mutual enhancement across tasks are demonstrated through multi-task transfer learning.
@@
  • Comprehensive Quantitative and Qualitative Experiments. EEGPT demonstrated superior performance across 12 datasets encompassing 5 tasks, surpassing both pretrain-then-finetune and training-from-scratch predominant specialist baselines. The effectiveness of our proposed method is further validated by extensive qualitative analyses.
2 Method
In this section, we elaborate on the comprehensive framework of EEGPT. A detailed framework is illustrated in Figure 3. We begin by representing a multi-electrode EEG signal as $x\in\mathbb{R}^{E_{i}\times T\times C}$, where $E_{i}$ represents the number of electrodes. For the entire signal, we segment it into $T$ one-second intervals. Each interval is represented by $C$ uniformly sampled points from the original signal.
2.1 Autoregressive Time Series Modeling
In this stage, we aim to develop a comprehensive and detailed self-supervised learning paradigm. It is designed to accurately and efficiently capture the intrinsic temporal variations in EEG signals. Furthermore, we plan for it to be electrode-conditioned, which will facilitate the discernment of both disparities and similarities across different electrodes, enhancing its effectiveness in diverse scenarios.
Electrode-wise Modeling Strategy.
We compile the pre-training dataset $X=\{x_{1},x_{2},\ldots,x_{N}\}$ by aggregating data from multiple sources. Each sample $x_{i}\in\mathbb{R}^{E_{i}\times T\times C}$ corresponds to a set $\mathcal{E}_{i}$ containing $E_{i}$ electrodes. To discern the distinctive patterns specific to each electrode in EEG signals, we introduce a structured reorganization function $\mathcal{R}(\cdot)$. Specifically, we segment each $x_{i}$ based on individual electrodes, denoted as $\mathbf{x}_{i}^{e}\in\mathbb{R}^{T\times C}$. Consequently, EEG recordings from disparate sources that share identical electrodes are grouped together:
$\mathcal{R}(X)=\left\{D_{e}\mid e\in\bigcup_{i=1}^{N}\mathcal{E}_{i}\right\}$ (1)
where $D_{e}$ is the grouped collection of all data segments $\mathbf{x}_{i}^{e}$ from electrode $e$ across all samples that include electrode $e$:
$D_{e}=\{\mathbf{x}_{i}^{e}\mid e\in\mathcal{E}_{i},\,i=1,2,\ldots,N\}$ (2)
@@ 2.1 Autoregressive Time Series Modeling
The size of $\mathcal{R}(X)$ corresponds to the count of unique electrodes present in $X$. To distinguish between different electrodes, we introduce a trainable electrode vocabulary $v_{E}\in\mathbb{R}^{|R(X)|\times C}$. All elements in $D_{e}$ share the same electrode embedding $v_{E}^{e}$. This embedding is served as condition and then concatenated along the sequence dimension to $x_{i}^{e}$. For simplicity, we consistently refer to the concatenated sequence as $x_{i}^{e}$:
$x_{i}^{e}=[v_{E}^{e}||x_{i}^{e}]\in\mathbb{R}^{(T+1)\times C}$ (3)
[…]
@@
Figure 3: Overview of the EEGPT architecture. (left) Autoregressive reconstruction serves as the pre-training objective, with a learnable condition token added at the start to distinguish between electrodes. Each signal token predicts the next token through the Electrode Temporal Encoder (ETE) one-by-one. (right) Electrodes in each dataset are processed through the pre-trained ETE, extracting the final token as the electrode representations, which are then fed into a Task-shared Electrode Graph (TEG) network to integrate spatial information across multiple electrodes. ETE and TEG collectively constitute a progressive spatiotemporal decoupling.
Autoregressive Reconstruction.
@@
As depicted in Figure 3 (left), each $x_{i}^{e}$ is inputted into a shared Electrode Temporal Encoder (ETE), which comprises a series of $L$ identical layers. Each layer contains two sub-layers: the first utilizes a multi-head causal attention mechanism, and the second employs a positionwise fully connected feed-forward network. Specifically, the input sequence $x_{i}^{e}$ first undergoes the causal attention process:
@@
where $Q$, $K$, and $V$ are queries, keys, and values respectively, all derived from $x_{i}^{e}$, and $d$ is the hidden size. $M$ is a causal mask designed to ensure that each token only attend to tokens that are sequentially prior to itself. The output of this sub-layer is then normalized and passed through a residual connection (He et al., 2016). Subsequently, it is fed into feed-forward network, which consists of two linear transformations with a SwiGLU (Dauphin et al., 2017) activation. Finally, the output from ETE is transformed into the corresponding next-token prediction through a simple MLP.
Training Objective.
Assuming the input is $x$, the reconstructed result is denoted as $\hat{x}$. Without loss of generality, the training objective for the autoregressive model can be formulated as follows:
$\mathcal{L}(\theta)=\frac{1}{T}\sum_{t=1}^{T}\rho(x_{t}-\hat{x}_{t}(x_{<t};\theta))$ (5)
where $\theta$ represents the model parameters that require optimization. The function $\rho(\cdot)$ serves as a distance metric to quantify the discrepancy between the actual and reconstructed values. For LLMs, due to the discrete nature of the vocabulary, Cross-Entropy (CE) is commonly used. Given the inherent continuity of EEG signals, we default to using Mean Squared Error (MSE) in this context.
2.2 Task-shared Electrode Graph

### v2

@@ Abstract
Electroencephalogram (EEG) signals are pivotal in providing insights into spontaneous brain activity, highlighting their significant importance in neuroscience research. However, the exploration of versatile EEG models is constrained by diverse data formats, outdated pre-training paradigms, and limited transfer learning methods, only leading to specialist models on single dataset or subject. In this paper, we introduce BrainGPT, the first generalist EEG foundation model designed to address these challenges. First, we propose an electrode-wise modeling strategy that treats the data of each electrode as an independent sample, enabling the integration of diverse EEG datasets collected from nearly all commonly used electrodes, amassing 37.5M pre-training samples. Second, we develop the first autoregressive EEG foundation model, moving away from traditional masked autoencoder approaches to a next-token-prediction task that better captures the temporal dependencies of EEG data. We also explore scaling laws with model up to billion level parameters — the largest scale in EEG research to date. Third, we introduce a multi-task transfer learning paradigm based on a learnable electrode graph network that is shared across tasks, which for the first time confirms multi-task compatibility and synergy. As a generalist EEG foundation model, BrainGPT shows broad compatibility with various signal acquisition devices, subjects, and tasks. It supports up to 138 electrodes and any combination thereof as input. Furthermore, we simultaneously evaluate it on 5 distinct downstream tasks across 12 benchmarks. BrainGPT consistently outperforms existing specialist models across all downstream tasks, with its effectiveness further validated through extensive ablation studies. This work sets a new direction for generalist EEG modeling, offering improved scalability and adaptability for a wide range of EEG applications. Both the training code and model checkpoints will be publicly available.
Index Terms:
EEG, Foundation Model, Autoregressive Pre-training, Generalist Model
I Introduction
@@
Electroencephalogram (EEG) captures spontaneous brain activity via electrograms, providing a non-invasive means to monitor neural dynamics [1]. Through the analyses of EEG, valuable insights are derived for various applications, including but not limited to emotion recognition [2], motor imagery classification [3], mental workload detection [4] and sleep stage classification [5]. This breadth of applications underscores the versatility and utility of EEG in neuroscientific research.
Figure 1: BrainGPT, as a generalist model, significantly outperforms dataset-specific specialist models across 12 benchmarks spanning 5 tasks. It strongly demonstrate the versatility and transferability.
Research on EEG downstream tasks has been thriving, yet most studies share a notable characteristic: specialization. At the data level, for instance, a variety of proprietary data formats [6, 7] and handcrafted feature extraction techniques [8, 9] have been developed to enhance the discriminability of domain-specific data. Similarly, at the model level, numerous specialized modules and architectures are designed and optimized for specific tasks [10], datasets [11], or even individual subjects [12]. While these approaches have advanced the field, they often lack generalizable and extensible capabilities. In contrast, an EEG model capable of simultaneously handling heterogeneous data and multiple tasks within a unified framework is highly anticipated. Such a model holds significant promise as a foundation for advancing EEG analysis by enhancing cross-domain adaptability, reducing the need for task-specific designs, and enabling more efficient learning.
Although some pioneering studies have made initial explorations into constructing EEG foundation models [13, 14, 15], three major challenges still remain:
@@
Data Format: EEG data exhibit significant heterogeneity [11, 16], characterized by variations in data collection devices and experimental tasks. Furthermore, different datasets may employ a diverse number and combination of electrodes based on practical considerations. The inconsistency in data formats across different sources hinders their integrations into a unified model for training, posing a significant challenge in the development of an EEG foundation model. Therefore, an efficient and scalable strategy for unifying these diverse EEG data format is extremely demanding.
Self-supervised Pre-training: Current EEG foundation models [14, 13, 15] predominantly rely on masked autoencoder (MAE) architectures [17, 18], which reconstruct masked portions of EEG signals using bidirectional attention mechanisms [19]. While effective in certain contexts, these approaches exhibit inherent limitations in modeling the temporal dependencies critical to EEG data. Specifically, MAE-based methods focus on reconstructing local signal segments, often neglecting the global temporal structure and long-range dependencies that are fundamental to understanding brain activity patterns [20, 21]. Considering that EEG signals often reflect a continuous and progressive flow of information, where past neural activity influences future states, the bidirectional nature of MAE architectures may not fully align this characteristic.
Transfer Learning. Current EEG foundation models are typically fine-tuned for specific datasets, resulting in specialists confined to narrow domains. However, this approach faces significant limitations in both efficiency and effectiveness. For efficiency, fine-tuning a separate model for each task is inherently costly, particularly given the increasing scale of models and the growing diversity of tasks. For effectiveness, such task-specific transfer fails to fully explore the potential for multi-task compatibility and synergistic interactions, thereby limiting the broader capabilities of foundation models.
@@
Based on the above considerations, we propose BrainGPT, the first generalist EEG foundation model offering extensive versatility as shown in Fig. 2. It seamlessly adapts and encodes signals collected by nearly all popular EEG acquisition devices, accommodating signals from up to 138 electrodes with various configurations and combinations. Moreover, it is capable of simultaneously supporting all prevalent downstream tasks within a single model, and it is highly scalable to new tasks. BrainGPT significantly diverges from previous methods, with its novelty encapsulated in three aspects:
@@
Figure 2: The versatility of BrainGPT is reflected in the broad compatibility with subjects, signal acquisition devices, and tasks. EEG signals from various subject, using various device, and performing various task can be characterized and analyzed effectively within one model.
1) For data format, we propose an electrode-wise modeling strategy, which deconstructs EEG signals by treating each electrode as an independent data sample. It enables the model to learn the unique patterns and temporal dynamics of each electrode, with a learnable special token assigned to each electrode as prefix condition. Based on it, we collect a total of 37.5M pre-training samples with approximately 1B tokens.
2) For the pre-training paradigm, we propose the first autoregressive EEG model, which effectively captures the temporal dependencies inherent in each electrode through the intuitive yet challenging task of next token prediction. We conduct pre-training across four scales (i.e., Base, Huge, Large, and Giant), with BrainGPT-Giant surpassing 1B parameters, making it the largest scale in the EEG field. Additionally, we firstly validate the scaling laws for both data and model size, confirming the scalability of this pre-training paradigm.
@@
3) For transfer learning, we propose a learnable graph network with electrodes as nodes. It is concurrently shared across multiple tasks. Node activation patterns are adaptively determined by corresponding input task. Leveraging the robust temporal representations learned from autoregressive pre-training, the electrode graph serves as a spatial supplement by integrating information from multiple electrodes. The whole framework is designed with a progressive spatiotemporal decoupling. We collect data from 12 benchmarks for joint learning instead of traditional single-task fine-tuning. These tasks achieves promising multi-task compatibility and synergism, making BrainGPT the first generalist EEG foundation model.
Our contributions are summarized as follows:
  • We introduce an electrode-wise modeling strategy, enabling flexible and scalable integration of heterogeneous EEG data. It allows BrainGPT to support up to 138 electrodes and arbitrary combinations, the largest coverage achieved so far, significantly enhancing generalizability.
  • We propose the first autoregressive pre-trained EEG model, which effectively captures the temporal dynamics inherent in EEG signals. We further validate the scaling laws of this pre-training paradigm with respect to data and model size, providing strong evidence for its scalability.
@@
  • Building upon the task-shared graph network, significant mutual enhancement across tasks are demonstrated, making BrainGPT the first generalist model to exhibit confirmed multi-task compatibility and synergism.
@@
  • BrainGPT demonstrates superior performance across 12 datasets encompassing 5 tasks, surpassing both pretrain-then-finetune and training-from-scratch predominant specialist baselines as shown in Fig. 1. The effectiveness of our proposed method is further validated by extensive qualitative analyses.
II Related Work
Despite recent exploratory studies on EEG foundation models gaining recognition, several challenges and limitations remain. Existing studies [14, 13, 15, 22, 23, 24] exhibit notable commonalities and limitations in data formatting strategies, self-supervised pre-training paradigm, and transfer learning, which may constrain their applicability and scalability.
Existing studies typically segment EEG signals into fixed-length tokens, which are either flattened into ”sentence”-like sequences or retain their spatiotemporal structure [14, 13, 23, 24]. To encode spatiotemporal information, fixed positional embeddings based on time or electrode location are commonly applied. While this formatting unifies diverse EEG inputs, it poses challenges. Notably, many methods only utilize subsets of electrodes during pretraining [14, 23], limiting scalability and generalization across datasets. The exclusion of signals from certain electrodes may impair performance on downstream tasks involving unseen channels. To address this, we propose an electrode-wise modeling strategy, treating each electrode’s signal as an independent data sample. This enables flexible electrode combinations and leverages full-scale EEG data (up to 138 electrodes) during pretraining.
In terms of self-supervised pre-training paradigm, most existing methods predominantly adopt MAE techniques. For instance, BIOT [14] and MMM [15] utilize channel and temporal embeddings to construct EEG tokens for masked segment prediction, whereas LaBraM [13] employs a neural tokenizer to discretize EEG signals and predict masked tokens based on visible patches. However, these MAE-based method may destroy the natural flow of brain signals, making them less suited to the intrinsic temporal dependencies of EEG signals. Hence, in this paper, we refine the current pre-training paradigms from bidirectional completion to sequential prediction. To be specific, we propose the first autoregressive-based self-supervised learning framework for EEG foundation model pre-training, aiming to more effectively capture the temporal structure inherent in EEG signals.
In terms of transfer learning, existing models predominantly rely on computationally intensive task-specific adaptation strategies, where pre-trained models are fine-tuned for each downstream task [14, 13, 22] or even individualized per subject [15]. This approach leads to significant inefficiencies, particularly given the prohibitive costs of repeated full parameter optimization [25], and fails to leverage cross-task synergies inherent in EEG data. To address these limitations, we propose a unified task-shared transfer learning paradigm, enabling BrainGPT to function as a generalist foundation model capable of simultaneous multi-task compatibility and synergistic knowledge transfer without task-specific retraining.
In summary, while self-supervised pre-training has shown promise in EEG analysis, existing approaches exhibit constraints in data formatting, self-supervised learning paradigms, and transfer learning methodologies. Addressing these limitations is crucial for enhancing the generalization and performance of EEG foundation models, and the proposed BrainGPT provides a promising solution.
III Method
In this section, we first introduce the electrode-wise modeling strategy adopted by BrainGPT, which decomposes the raw EEG signals on an electrode-wise basis to enhance flexibility. Subsequently, for each decomposed electrode signal, we propose a shared Electrode Temporal Encoder (ETE), which is pre-trained via autoregressive reconstruction to capture the temporal dynamics of individual electrodes. Finally, during the downstream task adaptation, we introduce the Task-shared Electrode Graph (TEG) network, designed to facilitate and enhance the collaborative learning of data across different tasks. The detailed framework is illustrated in Fig. 3.
III-A Electrode-wise Modeling Strategy
For a multi-electrode EEG signal, we first segment it into $T$ one-second intervals. Each interval is represented by $D$ uniformly sampled points from the original signal. We denote the segmented signal as $x_{i}\in\mathbb{R}^{E_{i}\times T\times D}$, where $E_{i}$ is the number of electrodes. Each EEG signal $x_{i}$ is associated with an electrode set $\mathcal{E}_{i}$ of size $E_{i}$. To facilitate electrode-wise modeling, we further decompose each $x_{i}$ into individual electrode signals $x_{i}^{e}\in\mathbb{R}^{T\times D}$. Each $x_{i}^{e}$ represents the temporal sequence of a single electrode and serves as a basic training sample. The complete pre-training dataset $X$ is then constructed as:
$X=\{\mathbf{x}_{i}^{e}\mid e\in\mathcal{E}_{i},\,i=1,2,\ldots,N\}$ (1)
The total number of electrodes covered in $X$ is given by:
$E=|\bigcup_{i=1}^{N}\mathcal{E}_{i}|.$ (2)
@@ III-A Electrode-wise Modeling Strategy
To differentiate between electrodes, we introduce a trainable electrode vocabulary $\mathcal{V}\in\mathbb{R}^{E\times D}$. All samples $x_{i}^{e}$ from the same electrode share the corresponding embedding $\mathcal{V}_{e}$. This embedding serves as a prefix condition and is concatenated along the sequence dimension with $x_{i}^{e}$. For simplicity, we continue to refer to the concatenated sequence as $x_{i}^{e}$.
$x_{i}^{e}=[\mathcal{V}_{e}||x_{i}^{e}]\in\mathbb{R}^{(T+1)\times D}$ (3)
where $\|$ signifies the concatenation operation. Hence, signals from various domains and electrodes have been standardized into a highly scalable format. The chronological sequences $x_{i}^{e}$, which contains T+1 EEG tokens, will serve as the fundamental unit for performing autoregressive reconstruction.
@@
Figure 3: Overview of the BrainGPT architecture. (left) Autoregressive reconstruction serves as the pre-training objective, with a learnable condition token added at the start to distinguish between electrodes. Each signal token predicts the next token through the Electrode Temporal Encoder (ETE) one-by-one. (right) Electrodes in each dataset are processed through the pre-trained ETE, extracting the final token as the electrode representations, which are then fed into a Task-shared Electrode Graph (TEG) network to integrate spatial information across multiple electrodes. ETE and TEG collectively constitute a progressive spatiotemporal decoupling.
III-B Autoregressive Pre-training
@@
As shown in Fig. 3 (left), each $x_{i}^{e}$ is fed into a shared Electrode Temporal Encoder (ETE), consisting of $L$ identical layers. Each layer comprises two sub-layers: a multi-head causal attention mechanism and a position-wise feed-forward network. Specifically, the input sequence $x_{i}^{e}$ first passes through the causal attention mechanism:
@@
$Q=W_{Q}x_{i}^{e},\quad K=W_{K}x_{i}^{e},\quad V=W_{V}x_{i}^{e}$ (5)
where $Q$, $K$, and $V$ are queries, keys, and values respectively, all derived from $x_{i}^{e}$ through corresponding learnable projection matrices $W_{Q}$, $W_{K}$ and $W_{V}$. The hidden size of ETE is denoted as $d$. $M$ is a causal mask designed to ensure that each token only attend to tokens that are sequentially prior to itself. The output of this sub-layer is then normalized and passed through a residual connection [26]. Subsequently, it is fed into feed-forward network, which consists of two linear transformations with a Swish [27] activation:
$\text{FFN}(x)=W_{\text{down}}\cdot\left(\text{Swish}(\mathbf{W}_{\text{gate}}\cdot x)\odot(\mathbf{W}_{\text{up}}\cdot x\right)$ (6)
Finally, we employ a lightweight MLP to predict the next token in the sequence based on the output of ETE. The model is trained by minimizing the reconstruction loss, conditioned on all preceding tokens:
$\mathcal{L}(\theta)=\frac{1}{T}\sum_{t=1}^{T}\rho(x^{e}_{i}[t]-\text{ETE}(x_{i}^{e}[\leq t];\theta))$ (7)
where $\theta$ denotes the trainable parameters of ETE. The function $\rho(\cdot)$ measures the discrepancy between the actual and reconstructed values, with Mean Squared Error (MSE) used as the default distance metric.
III-C Multi-task Transfer Learning